### v1

# TDS-CLIP: Temporal Difference Side Network for Image-to-Video Transfer Learning

Abstract
Recently, large-scale pre-trained vision-language models (e.g., CLIP), have garnered significant attention thanks to their powerful representative capabilities. This inspires researchers in transferring the knowledge from these large pre-trained models to other task-specific models, e.g., Video Action Recognition (VAR) models, via particularly leveraging side networks to enhance the efficiency of parameter-efficient fine-tuning (PEFT). However, current transferring approaches in VAR tend to directly transfer the frozen knowledge from large pre-trained models to action recognition networks with minimal cost, instead of exploiting the temporal modeling capabilities of the action recognition models themselves. Therefore, in this paper, we propose a memory-efficient Temporal Difference Side Network (TDS-CLIP) to balance knowledge transferring and temporal modeling, avoiding backpropagation in frozen parameter models. Specifically, we introduce a Temporal Difference Adapter (TD-Adapter), which can effectively capture local temporal differences in motion features to strengthen the model’s global temporal modeling capabilities. Furthermore, we designed a Side Motion Enhancement Adapter (SME-Adapter) to guide the proposed side network in efficiently learning the rich motion information in videos, thereby improving the side network’s ability to capture and learn motion information. Extensive experiments are conducted on three benchmark datasets, including Something-Something V1&V2, and Kinetics-400. Experimental results demonstrate that our approach achieves competitive performance. Code will be available at https://github.com/BBYL9413/TDS-CLIP.
In recent years, Large Language Models (LLMs) have achieved significant success (Mann et al. 2020; Touvron et al. 2023; Ryoo et al. 2021) in the fields such as text-video retrieval (Wu et al. 2023a), text-to-image generation (Qu et al. 2023), etc., gradually extending their influence to the domain of computer vision. Video Action Recognition (VAR) tasks, a challenging research area within computer vision, have seen mainstream large language models represented by the CLIP paradigm (Radford et al. 2021), which includes a Vision Transformer (ViT) vision encoder (Dosovitskiy et al. 2020) for image processing and a text encoder. CLIP establishes a strong association between images and text and is applied to various video tasks. This raises a thought-provoking question: how can we effectively utilize the pre-training of such large models for VAR tasks?
Figure 1: Comparison of per-video GFLOPs and accuracy on SSv2 (Goyal et al. 2017). The size of the bubbles represents the memory usage during training. EVL: Efficient Video Learning (Lin et al. 2022). ST-Adapter: Spatio-Temporal Adapter (Pan et al. 2022).
Recently, the emergence of Parameter-Efficient Fine-Tuning (PEFT) methods (Houlsby et al. 2019; Lester, Al-Rfou, and Constant 2021) has effectively alleviated this issue. PEFT adapts specific tasks by adjusting the learnability of certain parameters or introducing various adapters, avoiding extensive parameter adjustments across the entire model. With the advent of PEFT, researchers are exploring the possibility of freezing the original CLIP parameters and incorporating various adapters (Wang et al. 2024) or prompts (Wasim et al. 2023). These methods achieve impressive results in video understanding and are widely used in tasks such as video action recognition. As shown in Fig. 2(a), a popular design for efficient transfer learning is to insert tunable structures between pre-trained and parameter-frozen CLIP visual transformer blocks. Although introducing adapters under frozen parameters facilitates temporal action modeling, backpropagation still passes through the frozen layers, consuming memory. To enhance memory efficiency, LST (Sung, Cho, and Bansal 2022) introduces a side network connected to the frozen pre-trained model for backpropagation, as shown in Fig. 2(b). Continuing with this idea, Side4Video (Yao et al. 2023) further explores the capability of side networks in temporal modeling within the computer vision domain, achieving parameter-efficient video understanding tasks. These methods treat the spatial and temporal branches equally. However, for VAR tasks, it is often necessary to emphasize the effective processing of the temporal branch. The potential of using side networks for temporal modeling in video understanding tasks remains to be fully explored.
Figure 2: Comparison with commonly used and effective fine-tuning methods for video recognition. (a): ST-Adapter: Spatio-Temporal Adapter (Pan et al. 2022). (b): Ladder Side-tuning (Sung, Cho, and Bansal 2022).
In this work, we propose a method based on CLIP fine-tuning that balances memory efficiency and temporal modeling performance, called TDS-CLIP. First, we shift the adapter positions used for the CLIP vision branch to the side network to achieve memory-efficient pre-training fine-tuning. To better model the temporal information of videos, we design a new Temporal Difference Adapter (TD-Adapter) that effectively captures local temporal differences to enhance the model’s global temporal modeling capability. Second, we integrate the TD-Adapter with a multi-head self-attention module and feed-forward network to form a learnable side network module, used for parameter training and backpropagation. Finally, to improve the side network’s capability of capturing video motion information, we introduce a Side Motion Enhancement Adapter (SME-Adapter) that acts on the input of the side network to guide it in effectively learning additional motion information relevant to action labels.
Figure 3: (a) Illustration of the proposed TDS-CLIP for video understanding. TDS-CLIP consists of a frozen CLIP visual spatial encoder and a trainable visual-temporal encoder, where the visual temporal encoder is composed of the proposed SME-Adapter and TDS Block to form the side network, and the TDS Block is composed of the proposed TD-Adapter and ViT module. (b) Detailed Structure of the proposed TDS Block.
We evaluate the performance of our method on several large-scale temporal action recognition datasets, including Something-Something V1(Goyal et al. 2017), Something-Something V2(Goyal et al. 2017), and Kinetics-400(Kay et al. 2017). Experimental results demonstrate that the proposed method effectively accomplishes video action understanding tasks.
Our contributions are summarized as follows:
  • We propose a novel memory-efficient Temporal Difference Side Network (TDS-CLIP) for video action recognition that balances knowledge transfer and temporal modeling without requiring backpropagation in models with frozen parameters
  • We propose the Temporal Difference Adapter (TD-Adapter) and the Side Motion Enhancement Adapter (SME-Adapter), which enhance the model’s ability to capture and learn from motion information, thereby improving both local and global temporal analysis.
  • Extensive experiments evaluate the efficacy of TDS-CLIP on public datasets, demonstrating competitive performance against state-of-the-art approaches.
Related Works
VLM for Video Recognition. ViT (Dosovitskiy et al. 2020) models are gradually replacing convolutional neural networks in the mainstream due to their powerful learning capabilities and flexibility. EVA-CLIP (Sun et al. 2023) proposes a 64-layer ViT-E with larger scale parameters, and the effective adaptation of such a huge image model to the video domain is a task well worth exploring. Meanwhile, Transferring pre-trained models of language-images from the ViT architecture to video attracts widespread attention (Cheng et al. 2021). Previous action recognition methods rely on pre-trained models from ImageNet (Simonyan and Zisserman 2014). Currently, powerful text-image pre-trained models based on CLIP are gradually emerging and are increasingly used in VAR tasks (Wang et al. 2023). EVL (Lin et al. 2022) uses parallel Transformer decoders to extract spatial features from the frozen CLIP model for VAR tasks. ST-Adapter (Pan et al. 2022) introduces a spatiotemporal convolution adapter to endow the CLIP model with the capability of understanding videos. AIM (Yang et al. 2023) introduces spatial adaptation, temporal adaptation, and joint adaptation, gradually equipping image models with spatiotemporal reasoning abilities, further enriching the types of adapters to suit different video action recognition tasks. DUALPATH (Park et al. 2023) achieves dynamic modeling using spatial and temporal adaptation paths, merging successive frames into a grid-like frameset to accurately infer relationships between markers. M2-CLIP (Wang et al. 2024) introduces a TED-Adapter, enhancing the model’s generalization ability while maintaining video action recognition performance, and successfully uses it for multimodal VAR tasks. These methods effectively transfer the powerful pre-trained CLIP visual encoder to VAR tasks but do not address the issue of memory consumption during backpropagation.
Side Network. The side-tuning network (Zhang et al. 2020) is first proposed to address overfitting and the problem of forgetting in incremental learning, and it applies to various downstream tasks. In vision tasks, LST (Sung, Cho, and Bansal 2022) tackles this problem by implementing a side branch network attached to the CLIP visual model. This scheme is effective in avoiding the passing of gradients in the visual encoder of frozen parameters thus reducing memory usage. Inspired by the LST, DiST (Qing et al. 2023) uses a dual encoder structure, employing the pre-trained CLIP base model as the spatial encoder and introducing a lightweight additional network as the temporal encoder, finally integrating spatiotemporal information through an ensemble branch to achieve memory-efficient video understanding. Side4Video (Yao et al. 2023) further explores the effective forms of memory-efficient side networks in video action recognition tasks and proposes a spatiotemporal side network for video understanding tasks. Although these methods demonstrate that using side networks can achieve memory-efficient image-to-video transfer learning, they do not deeply explore the temporal modeling capabilities of side network models, which we further enhance. In this work, we introduce corresponding temporal modeling adapters at both the input and the path of the side network composed of ViT encoders, transforming it into a memory-efficient and powerful temporal modeling framework for VAR tasks.
Method
As shown in Fig. 3, the overall architecture of TDS-CLIP framework consists of two key components: the visual spatial encoder and the visual temporal encoder. The visual spatial encoder is composed of a frozen CLIP pre-trained ViT, which provides strong spatial semantic features for the input video frames. The visual temporal encoder is a side network structure that accepts the output features from the visual spatial encoder. It takes the dense video frame features provided by the SME-Adapter as input and transforms the spatial features from the frozen CLIP branch into an effective form suitable for temporal modeling in video understanding.
Here, we present an overview of the entire architecture and present the details of the visual spatial encoder and the visual temporal encoder in the following two sections. The input video frames to the model are formally defined as $V\in\mathbb{R}^{3\times T\times H\times W}$, where $H$ and $W$ represent spatial dimensions, and $T$ represents the sampled frames.
Figure 4: Detailed structure of the proposed SME-Adapter. We illustrate the feature map as the original RGB frame for better intuitive understanding. “Diff” indicates difference operation.
Visual Spatial Encoder
The spatial encoder consists of the ViT with frozen parameters from the CLIP visual branch, which can independently capture the spatial semantic features of sparse input video frames, thereby leveraging its powerful pre-training to enhance training performance. Following the ViT format, the input frames are first divided into $N=\frac{HW}{P^{2}}$ patches, each with a shape of $P^{2}$, and then projected through a fully connected layer. Then, for the input full-length video frame sequence, each patch feature vector is represented as $X_{i,j}$, where $i=\{1,2,...,T\}$ and $j=\{1,2,...,N\}$ denotes the patch index. Additionally, a learnable $E_{cls}$ token and a trainable position embedding $E_{pos}$ are added to each frame. With a total of $L$ transformer blocks, the input format can be represented as:
$Z^{l}=Transformer^{l}([E_{cls};{X_{i,j}^{l-1}]+E_{pos}}),$ (1)
where $l=\{1,2,...,L\}$ denotes the number of layers of the ViT, The output shape is $Z^{l}\in\mathbb{R}^{T\times(N+1)\times C}$.
Method Pre-train Frames×$\times$×Cr.$\times$×Cl. TFLOPs SSv1(%) SSv2(%)
Top-1 Top-5 Top-1 Top-5
Full Fine-tuning
ViViT-L(Arnab et al. 2021) IN-21K 32$\times$3$\times$4 1.0$\times$12 - - 65.9 89.9
Video Swim-B(Liu et al. 2022b) IN-21K 32$\times$3$\times$1 0.3$\times$3 - - 69.6 92.7
MViT-B(Fan et al. 2021) Kinetics-600 32$\times$3$\times$1 0.2$\times$3 - - 68.7 91.5
UniFormer V2-L/14(Li et al. 2023) CLIP-400M 16$\times$3$\times$1 0.9$\times$3 60.5 86.5 72.1 93.6
ILA-ViT-L/14(Tu et al. 2023) CLIP-400M 16$\times$3$\times$4 0.9$\times$12 - - 67.8 90.5
S-ViT-B/16(Zhao et al. 2023) CLIP-400M 32$\times$3$\times$2 0.3$\times$6 - - 69.3 92.1
Frozen CLIP
EVL-L/14(Lin et al. 2022) CLIP-400M 8$\times$3$\times$1 0.8$\times$3 - - 65.1 -
ST-Adapter-L/14(Pan et al. 2022) CLIP-400M 16$\times$3$\times$1 1.4$\times$3 - - 71.9 93.4
DUALPATH-L/14(Park et al. 2023) CLIP-400M 16$\times$3$\times$1 0.5$\times$3 - - 70.2 92.7
AIM-L/14(Yang et al. 2023) CLIP-400M 16$\times$3$\times$1 1.9$\times$3 - - 69.4 92.3
DiST-L/14(Qing et al. 2023) CLIP-400M 16$\times$3$\times$1 1.4$\times$3 - - 72.5 93.0
M2-CLIP-B/16(Wang et al. 2024) CLIP-400M 32$\times$3$\times$1 0.8$\times$3 - - 69.1 -
Side4Video-B/16(Yao et al. 2023) CLIP-400M 16$\times$3$\times$2 0.4$\times$6 60.7 86.0 71.5 92.8
Ours (Frozen CLIP)
TDS-CLIP-B/16 CLIP-400M 8$\times$3$\times$2 0.2$\times$6 60.1 85.4 71.8 93.0
TDS-CLIP-B/16 CLIP-400M 16$\times$3$\times$2 0.4$\times$6 61.0 86.3 72.1 93.3
TDS-CLIP-L/14 CLIP-400M 8$\times$3$\times$2 0.8$\times$6 63.0 87.8 73.4 93.8
Table 1: Comparison of performance with state-of-the-art methods on SSv1 & SSv2. “Cr.” and “Cl.” are the abbreviation for “spatial crops” and “temporal clips”.
Visual Temporal Encoder
The spatial semantic features from the frozen CLIP ViT can guide the model in learning the appearance and background information of video frames, but for video understanding tasks, the model needs to have temporal modeling capabilities. The visual temporal encoder is designed as a trainable side network structure that accepts the output from the visual spatial encoder, featuring its own independent video feature input mode and trainable ViT modules. Specifically, for the video frames input, we aim to guide the side network to effectively learn the latent motion information in the video through a Side Motion Enhancement Adapter (SME-Adapter). For the trainable ViT module, we design a Temporal Difference Adapter (TD-Adapter) to integrate the CLIP spatial semantic information with the temporal information from the side network, thereby assisting the model in performing effective temporal modeling.
SME-Adapter:
Specifically, our SME-Adapter performs low-level feature extraction at the beginning of the side network, utilizing temporal differences between local RGB frames to model local motion features. As shown in Fig.  4, for the input video sequence $V\in\mathbb{R}^{3\times T\times H\times W}$, assuming that the i-frame is $V_{i}$, we first extend the sampling by $n$ frames centered on $V_{i}$, formally denoted as $V_{i}^{n}=\{V_{i-n},...,V_{i},...,V_{i+n}\}$. Then, we extract the feature differences $D_{i}$ between frames within a local sampling window centered on $V_{i}^{n}$, and finally concatenate them along the channel dimension. This effective form is proposed as:
$D_{i}=Concat(V_{i-n+1}-V_{i-n},...,V_{i+n}-V_{i+n-1})$ (2)
The extracted features are processed through a 3D convolution module to align with the shape of the original features after convolution. Subsequently, feature fusion is performed via element-wise addition, resulting in the final SME-Adapter output features $S_{out}^{i}$. Formally, this stage can be represented as:
$S_{out}^{i}=\alpha Conv1(D_{i})+\beta Conv2(V_{i}),$ (3)
where, the $\alpha$ and $\beta$ represent adjustment coefficients. $S_{out}^{i}$ will be used as an input to the side network and integrated with the VIT output of the first frozen CLIP into the TDS Block.
Method Pre-train Frames×$\times$×Cr.$\times$×Cl. TFLOPs Top-1(%) Top-5(%)
Full Fine-tuning
Video Swim-L(Liu et al. 2022b) IN-21K 32$\times$3$\times$4 0.6$\times$12 83.1 95.9
ViViT FE-L(Arnab et al. 2021) IN-21K 128$\times$3$\times$1 4.0$\times$3 81.7 93.8
MViT-B(Fan et al. 2021) Kinetics-600 32$\times$5$\times$1 0.2$\times$5 82.9 95.7
Uniformer V2-B/16(Li et al. 2023) CLIP-400M 8$\times$3$\times$4 0.2$\times$12 85.6 97.0
BIKE-L/14(Wu et al. 2023b) CLIP-400M 16$\times$3$\times$4 0.8$\times$12 88.1 97.9
ActionCLIP-B/16(Wang et al. 2023) CLIP-400M 32$\times$3$\times$10 0.6$\times$30 83.8 96.2
X-CLIP-B/16(Wang et al. 2023) CLIP-400M 16$\times$3$\times$4 0.6$\times$12 84.7 96.8
S-ViT-B/16(Zhao et al. 2023) CLIP-400M 16$\times$3$\times$4 0.7$\times$12 84.7 96.8
Frozen CLIP
EVL-B/16(Lin et al. 2022) CLIP-400M 8$\times$3$\times$1 0.4$\times$3 82.9 -
ST-Adapter-B/16(Pan et al. 2022) CLIP-400M 8$\times$3$\times$1 0.2$\times$3 82.0 95.7
ST-Adapter-B/16(Pan et al. 2022) CLIP-400M 32$\times$3$\times$1 0.6$\times$3 82.7 96.2
DUALPATH-B/16(Park et al. 2023) CLIP-400M 32$\times$3$\times$1 0.7$\times$3 85.4 97.1
AIM-B/16(Yang et al. 2023) CLIP-400M 8$\times$3$\times$1 0.2$\times$3 83.9 96.3
AIM-B/16(Yang et al. 2023) CLIP-400M 32$\times$3$\times$1 0.6$\times$3 84.7 96.7
DiST-B/16(Qing et al. 2023) CLIP-400M 8$\times$3$\times$1 0.2$\times$3 83.6 96.3
DiST-B/16(Qing et al. 2023) CLIP-400M 32$\times$3$\times$1 0.7$\times$3 85.0 97.0
M2-CLIP-B/16(Wang et al. 2024) CLIP-400M 32$\times$3$\times$4 0.8$\times$12 84.1 96.8
Side4Video-B/16(Yao et al. 2023) CLIP-400M 32$\times$3$\times$4 0.7$\times$12 84.2 96.5
Ours (Frozen CLIP)
TDS-CLIP-B/16 CLIP-400M 8$\times$3$\times$4 0.2$\times$12 83.9 96.1
TDS-CLIP-B/16 CLIP-400M 32$\times$3$\times$4 0.7$\times$12 84.5 96.7
TDS-CLIP-L/14 CLIP-400M 8$\times$3$\times$4 0.8$\times$12 86.7 97.5
Table 2: Comparison of performance with state-of-the-art methods on Kinetics-400. “Cr.” and “Cl.” are the abbreviation for “spatial crops” and “temporal clips”.
TD-Adapter:
To transfer the learning capability of the CLIP visual branch to video tasks, an additional temporal modeling module is required. Thus, we design a TD-Adapter in the side network, which accepts both spatial features from the frozen ViT and temporal input features from the side network, transforming them into encoder features suitable for video temporal modeling.
As shown in Fig. 3(b), the input feature of the TD-Adapter is denoted as $Z_{v}\in\mathbb{R}^{T\times N\times C}$. First, it is reshaped to a form $Z_{s}\in\mathbb{R}^{C\times T\times H\times W}$ suitable for 3D convolution. Then, a set of convolution operations is used for dimensionality reduction to obtain the feature $Z_{s}\in\mathbb{R}^{C/r\times T\times H\times W}$, $r$ denotes rate of reduction. Next, we capture the sensitive feature values of the local temporal dimension features using a 3D max-pooling operation, and a feature difference operation is employed to obtain dynamic features of local motion patterns.
$D_{z}=Z_{s}-Pool(Z_{s}),$ (4)
where $D_{z}$ denotes the feature difference. This method is commonly used in the early stages of convolutional models for efficient temporal modeling (Wang, She, and Smolic 2021; Li et al. 2020; Wang et al. 2021).
Finally, the reshaping operation and a set of convolution operations are again used to recover the feature shapes suitable for the ViT encoder.
TDS-CLIP for video action recognition:
We integrate the proposed SME-Adapter and TD-Adapter into Side4Video(Yao et al. 2023) architecture, where the TD-Adapter, combined with a trainable ViT encoder, forms the TDS-CLIP block. We retain the CLS token shift(Liu et al. 2022a; Zhang, Hao, and Ngo 2021) operation to further improve the capacity for learning temporal information, offering the advantage of minimal additional memory overhead. The number of these blocks matches the frozen ViT encoders in the visual branch of CLIP and combined via side-connections. For action recognition tasks, we obtain the prediction scores by performing global average pooling on the output of the final TDS-CLIP block.
Experiments
Experimental Setup
To evaluate the effectiveness of our method, we assess it on several widely used action recognition benchmark datasets, including Something-Something V1 (SSv1) (Goyal et al. 2017), Something-Something V2 (SSv2) (Goyal et al. 2017), and Kinetics400 (Kay et al. 2017). SSv1 and SSv2 are commonly used datasets for evaluating the importance of temporal modeling. Kinetics400 is a comprehensive action recognition dataset spanning 400 different human actions. We use the CLIP pretrained ViT-B/16 and ViT-L/14 as spatial encoders. By adjusting the dimensions, our model balances memory usage and performance. Due to hardware dilemmas, unless otherwise specified, we use the corresponding ViT-B/16 or ViT-L/14 structure in the temporal encoder to form the TDS-CLIP framework, employing a sparse frame sampling strategy of 8 or 16 frames during both training and inference.
Component Top-1(%) Acc.
Baseline 58.1
+ TD-Adapter 59.1
+ SME-Adapter 59.4
(a) Study of component effectiveness.
Number Top-1(%) Acc.
1 58.0
3 58.5
5 59.4
(b) The impact of local frame sample number.
$\alpha$ / $\beta$ Top-1(%) Acc.
0.5 58.4
1.0 59.4
1.5 58.9
(c) The impact of fusion rate.
Method Top-1(%) Acc.
Pool 59.4
Conv 59.3
Direct 57.3
(d) The impact of temporal difference module.
Kernel Size Top-1(%) Acc.
k=3 59.4
k=5 59.2
k=7 58.8
(e) The impact of kernel size in pooling layers.
Embedding Top-1(%) Acc.
128 57.4
320 59.4
512 58.2
(f) The impact of dimensionality.
Table 3: Ablation study results on Something-Something V1. Unless otherwise specified, all models use ViT-B/16 with 8 frames under the single view protocol.
Comparison with State-of-the-art
Results on Something-Something V1&V2:
Tab. 1 shows the comparison results with other state-of-the-art methods on SSv1 and SSv2. Both fully fine-tuning and frozen CLIP parameter methods, our approach achieves competitive results on both SSv1 and SSv2. For instance, in the fully fine-tuning approach, UniFormer-V2-L/14(Li et al. 2023) achieves the best classification results with the ViT-L/14 backbone, while our TDS-CLIP demonstrates superior performance with the same backbone (63.0% vs. 60.5% on SSv1; 73.4% vs. 72.1% on SSv2) while using fewer frames. In the frozen CLIP parameter methods, Side4Video-B/16(Yao et al. 2023) uses the same side network structure; in contrast, TDS-CLIP achieves better results with the same frame and sampling settings (61.0% vs. 60.7% on SSv1; 72.1% vs. 71.5% on SSv2), indicating that the proposed adapter enables effective temporal modelling. Additionally, it can be observed that the ViT-L/14 backbone outperforms the ViT-B/16 backbone under the same settings. These experimental results show that the proposed TD-Adapter and SME-Adapter are effective methods for enhancing temporal modeling and facilitating the transfer of knowledge from images to videos in frozen CLIP models.
Results on Kinetics-400:
To evaluate the generalization performance of our method, we conduct experiments on the K400 dataset. Compared to SSv1 and SSv2, the K400 emphasizes learning appearance features. As shown in Tab. 2, our method underperforms BIKE-L/14 (Wu et al. 2023b) in terms of appearance feature learning capability when compared to fully fine-tuned approaches. However, with 8-frame inputs, our performance surpasses that of ActionCLIP-B/16 (Wang et al. 2023), MViT-B (Fan et al. 2021), and Video Swin-L (Liu et al. 2022b) with 32-frame inputs, while also maintaining lower computational cost. When compared to methods that freeze CLIP, our performance with 32-frame inputs is lower than DiST-B/16 (Qing et al. 2023), AIM-B/16 (Yang et al. 2023), and DUALPATH-B/16 (Park et al. 2023), which possess independent spatial feature learning mechanisms. In contrast, TDS-CLIP focuses more on temporal feature learning and modeling. Notably, we achieve a competitive advantage when compared to methods like DiST-B/16 (Qing et al. 2023), AIM-B/16 (Yang et al. 2023), ST-Adapter-B/16 (Pan et al. 2022), and EVL-B/16 (Lin et al. 2022) with 8-frame inputs, indicating that TDS-CLIP captures more comprehensive video action information with fewer input frames. Lastly, we also evaluate the performance of TDS-CLIP with an 8-frame ViT-L framework, demonstrating that larger-scale frameworks are effective in improving performance. The experimental results demonstrate that TDS-CLIP not only excels in datasets with strong temporal correlations but also performs well in datasets with significant appearance background correlations. With these observations, we can conclude that TDS-CLIP has the dual advantage of spatial modeling and temporal modeling, showcasing robust generalization capabilities.
Ablation and Analysis
In this section, unless otherwise specifed, we use ViT-B/16 as the backbone and the 8 input frames on Something-Something V1.
Study of component effectiveness.
We explore the effectiveness of the proposed adapter and the results are shown in Tab. 3(a). The conventional 3D Conv is used to replace the TD-Adapter for the baseline of comparison. We can observe that the classification accuracy is improved by 1.0% after adding TD-Adapter, which proves that TD-Adapter can effectively improve the temporal modelling capability. Then, we add SME-Adapter, and the accuracy is improved even further, which implies that the SME-Adapter guides the side network to learn the motion features to further enhance the temporal modelling capability of the model, and proves that TD-Adapter and SME-Adapter are complementary to each other in terms of performance improvement.
The impact of local frame sample number.
We compare the performance impact of different numbers of local sampling video frames, and the results are shown in Tab. 3(b). 1 indicates that the SME-Adapter is not enabled, and serves as the baseline for comparison. The accuracy improves by 0.5% when the number is increased to 3 and by 1.4% when the number is increased to 5. This shows that increasing the number of local sampling frames can effectively improve the recognition performance. We do not test larger quantities because the total number of video frames used for testing is not large and bigger quantities may destroy the sparse sampling efficiency.
The impact of fusion rate.
The fusion ratio represents the proportion of the proposed method in feature fusion. The test results, as shown in Tab. 3(c), indicate that performance is optimal under equal-ratio fusion, while the performance deteriorates when the output feature ratio of the SME-Adapter is either too high or too low. This shows that the appropriate equal-ratio feature fusion method can well exploit the ability of SME-Adapter to learn and capture motion features.
Ablations for TD-Adapter.
We explore the impact of different core modules of the TD-Adapter on performance, including temporal convolutional layer enhancement(Wang et al. 2024) and direct feature difference calculation, as shown in Tab. 3(d). Using pooling layers achieves a classification accuracy of 59.4%, while temporal convolutional layer enhancement yields a slightly lower accuracy, decreasing by 0.2%. Direct feature difference calculation yields the lowest accuracy, as this approach tends to cause the loss of spatial features. It is noteworthy that pooling is a parameter-free operation, which helps reduce the number of parameters.
Furthermore, we test different kernel sizes for the pooling layers, specifically k=3, 5, and 7, as illustrated in Tab. 3(e). The classification accuracy decreases as the kernel size increases, indicating that an excessively large kernel size may disrupt the temporal relationships between features, therefore, larger kernel sizes should not be selected.
The impact of dimensionality.
In the transformer architecture, adjusting the dimensionality can effectively control the model’s memory consumption and impact its performance. Generally, increasing the dimensionality enhances the model’s complexity and memory usage, while also improving the modeling capacity of different layers. We follow the baseline comparison methods to test the impact of various dimensionality settings on performance, as shown in Tab. 3(f). The results show that the model achieves the best classification performance at 320 dimensions, while both excessively large or small dimensions constrain performance.
Layer 1-6 Layer 7-12 SME-Adapter Top-1(%) Acc.
58.8
58.6
59.1
59.4
Table 4: Study on the location of TD-Adapter layers.
Study on the location of TD-Adapter layers.
We study the location of the TD-Adapter within the structure, with the results shown in Tab. 4. We experiment with adding the TD-Adapter to the first half of the network (layers 1-6) and the second half (layers 7-12). Replacement with normal 3D convolution when not using TD-Adapter. The table shows that while using more TD-Adapters generally leads to better classification results, adding them to shallower layers provides greater benefits than those obtained from deeper layers. Additionally, the performance is further improved when the SME-Adapter is also used.
Figure 5: Visualisation of activation maps for features. Blue colour indicates larger magnitude of feature activation values and red colour indicates smaller magnitude of feature activation values.
Visualisation Related.
As show in Fig. 5, we further visualize the activation maps of features in the side network with and without using the SME-Adapter. It is evident that the SME-Adapter is more sensitive to the motion areas of the target and does not show a bias towards background information. For example, our SME-Adapter focuses more on the motion region of the hand that interacts with the object, rather than all the targets such as flower pots, water bottles, or mobile phones that appear in the scene, which is more in line with the human perception. This indicates that the proposed SME-Adapter facilitates the learning of motion features in videos, enabling the model to focus more accurately on dynamically moving target objects.
Conclusion
In this paper, we introduce a novel approach to VAR task that balances memory efficiency with effective temporal modeling. Our method addresses the challenging task of transferring large, task-specific vision-language models to the domain of video action recognition, which can be facilitated for researchers with limited hardware. The core innovation of our work is the development of a temporal modeling adapter tailored for side networks and the integration of spatio-temporal features, all while circumventing the need for backpropagation through the CLIP pre-trained model. Comprehensive experiments conducted on various benchmark datasets demonstrate that our method significantly enhances temporal modeling capabilities. In the era of large language models dominated by Transformers structures, we hope our work will inspire further exploration of parameter-efficient fine-tuning methods for video temporal modeling.